OpenStudy (anonymous):

Medal and a fan!!! PLZ HELP! Given a polynomial f(x), if (x + 7) is a factor, what else must be true?

OpenStudy (anonymous):

f(0) = 7 f(0) = -7 f(-7) = 0 f(7) = 0

OpenStudy (anonymous):

kk i can help you

OpenStudy (anonymous):

we know that what ever we get it has to be added with 7.

OpenStudy (anonymous):

if you picked D you get f(7) but once you add that 7 with 7 it will not equal 0.

OpenStudy (anonymous):

if you pick B 0+7 will be 7 not -7 so B isn't it either.

OpenStudy (anonymous):

so the answer is c?

OpenStudy (anonymous):

what confuses me is A and C are both true...

OpenStudy (anonymous):

but i didnt figure it out...

OpenStudy (anonymous):

its ok

OpenStudy (mathmate):

@vineeth10 It is a known property of polynomials that when a polynomial p(x) is divided by (x-k), the remainder equals p(k).

OpenStudy (mathmate):

For example, we know that \(p(x)=(x^2+3x+2)=(x+1)(x+2)\), so (x+1) and (x+2) are factors of p(x). This in turn means that p(-1)=0, and p(-2)=0.

OpenStudy (anonymous):

@mathmate or anyone The options are:f(0) = 7 f(0) = -7 f(-7) = 0 f(7) = 0

OpenStudy (aum):

If (x + 7) is a factor of f(x) it means, f(x) can be written as follows: f(x) = (x+7) * (some other factor) If I put x = -7 above, I will get: f(-7) = (-7 + 7) * (some other factor) f(-7) = 0 * (some other factor) f(-7) = 0

OpenStudy (anonymous):

Thanks!!! amu

OpenStudy (aum):

yw.